Circular No. 7253

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

GM SAGITTARII AND SAX J1819.3–2525 = XTE J1819–254 R. Stubbings, Drouin, Vic., Australia, reports that GM Sgr is undergo-

ing a very bright outburst after showing recent rapid variations, his visual magnitude estimates being: Aug. 11.488 UT, 12.3; 14.436, 12.5; 29.424, 12.7; 31.496, 13.0; Sept. 9.462, 11.2; 11.599, 11.9; 12.409, 11.1; 13.397, 11.9; 14.491, 11.4; 15.395, 8.8; 15.408, 8.8. Additional m_{v} estimates by A. Pearce, Nedlands, W. Australia: Sept. 15.550, 9.5; 15.571, 9.6.

D. A. Smith, A. M. Levine, and E. H. Morgan, Massachusetts Institute of Technology (MIT), on behalf of the RXTE All-Sky Monitor (ASM) team at MIT and Goddard Space Flight Center, report that, in response to the visual observations of GM Sgr, SAX J1819.3–2525 (an x-ray transient likely to be associated with GM Sgr; cf. IAUC 7119, 7120) was measured by the ASM on Sept. 15.420 UT to have an intensity of 1.55 ± 0.04 Crab (2– 12 keV; estimated error based solely on counting statistics). Subsequent measurements showed a rapid increase in x-ray intensity: Sept. 15.422, 1.32 ± 0.04 Crab; 15.423, 1.43 ± 0.04; 15.488, 2.41 ± 0.04; 15.489, 3.74 ± 0.05; 15.701, 10.0 ± 0.2; 15.703, 12.2 ± 0.2. The spectrum is harder than that of the Crab nebula.

The star attributed by Goranskij (1990, IBVS 3464) to GM Sgr has been measured on the Digital Sky Survey by G. V. Williams, Smithsonian Astrophysical Observatory, to be at α = 18^{h}19^{m}21^{s}.61, δ = −25^{o}24′26′′.3 (equinox 2000.0; uncertainty ± 1′′.0 in both coordinates); this is 1′ north of the GCVS position from Luyten (1927, HCO Bull. No. 852).

COMET P/1999 RO_{28} (LONEOS) C. Hergenrother, Lunar and Planetary Laboratory, reports that a co-

added 600-s CCD exposure with the Steward Observatory 2.3-m reflector on Sept. 13 of 1999 RO_{28} (discovered by LONEOS; cf. MPEC 1999-R23; discovery position given below) shows a stellar condensation with a faint 20′′ tail in p.a. 310^{o}. M. Tichý and J. Tichá, Kleť, later reported a faint coma of diameter 8′′ and 7′′ on images taken on Sept. 8.93 and 10.02 UT, respectively. Observations by J. V. Scotti with the Spacewatch telescope at Kitt Peak on Sept. 15.4 show a coma diameter of 9′′ (m_{1} = 18.6–18.7, m_{2} = 20.6–21.0) and a 0′.72 tail in p.a. 308^{o}. New orbital elements on MPEC 1999-R47 include T = 1999 Oct. 2.35 TT, q = 1.2302 AU, P = 6.46 yr.

1999 UT  α_{2000}  δ_{2000}  m_{2}
Sept.7.32994  23^{h}41^{m}39.^{s}78  −14^{o}32′00′′.5  18.4
1999 September 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green